Circular No. 8611

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ds, 2005em, 2005er, 2005es, 2005et, 2005eu Further to IAUC 8608, W. Li reports the LOSS discovery of an ap-

parent supernova on unfiltered KAIT images taken on Oct. 4.37 (at mag 17.9) and 5.36 UT (mag 17.6). SN 2005eu is located at α = 2^{h}27^{m}43^{s}.26, δ = +28^{o}10′36′′.6 (equinox 2000.0), which is 0′′.9 west and 0′′.8 south of the nucleus of the a small, faint galaxy that is near IC 226. A KAIT image taken on Sept. 15.35 showed nothing at this position (limiting mag ∼ 19.5).

N. Morrell, G. Folatelli, and M. M. Phillips, on behalf of the Carnegie Supernova Project, report spectroscopic observations (range 380–920 nm) of SN 2005er (cf. IAUC 8608), SN 2005es (cf. IAUC 8608), SN 2005et (cf. IAUC 8610), and SN 2005em (cf. IAUC 8604) obtained on Oct. 5.13, 5.18, 5.27, and 5.35 UT, respectively, with the Las Campanas 2.5-m du Pont telescope (+ WFCCD spectrograph). SN 2005em is probably a type-IIb supernova a few days after maximum light, the spectrogram showing some resemblance to those of SN 1987K. SN 2005er is a peculiar type-Ia supernova, similar to SN 1991bg, probably a few days after maximum light. Strong Ti II absorptions are observed. Assuming the NED recession velocity of 7847 km/s for the host galaxy (from Wegner et al. 1999, MNRAS 305, 259), an expansion velocity of 9600 km/s is derived from the absorption minimum of the Si II 635.5-nm doublet. SN 2005es is a young type-II supernova showing a rather featureless, blue continuum in the somewhatnoisy spectrogram. Hα and Hβ are seen with P-Cyg profiles. Assuming the NED recession velocity of 11287 km/s for the host galaxy (from Huchra et al. 1999, Ap.J. Suppl. 121, 287), an expansion velocity of 9600 km/s is derived from the absorption minimum of Hβ. SN 2005et is a type-Ia supernova about one week after maximum light. The spectrogram is similar to that of SN 1989B at eight days after maximum; an expansion velocity of 10000 km/s is derived from the minimum of the Si II 635.5-nm line, adopting the NED recession velocity of 10371 km/s for the host galaxy (from Wegner et al., op.cit.).

R. J. Foley and A. V. Filippenko, University of California at Berkeley; and T. Matheson, National Optical Astronomy Observatory, report that inspection of CCD spectra (range 470–990 nm), obtained on Oct. 5 UT with the Keck II 10-m telescope (+ DEIMOS), shows that SN 2005er is of type Ia, resembling SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543) near maximum brightness. Spectra of SN 2005es show that it is a young type-II supernova. Spectra of SN 2005ds (IAUC 8592) confirm that it is of type IIn, as suspected previously (IAUC 8594).

2005 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green